# 111preprint

### v1

@@ 6. Discussion
The concept of trust has been explored across various domains, including both human-human and human-robot interactions. Mayer et. al. (mayer1995integrative, ) formally define trust between individuals as “the willingness of a party to be vulnerable to the outcomes of another party based on the expectation that the other will perform a particular action important to the trustor, irrespective of the ability to monitor or control that other party.”.
The definition emphasizes that trust becomes crucial when vulnerability, risk, or uncertainty is involved in an interaction. With the deployment of robots working in close proximity with human partners in risky and uncertain environments (takayama2008beyond, ), human trust in robots becomes a crucial factor for their adoption. While human-robot trust shares commonality with human-human trust, the former is typically considered only from the human’s perspective regarding the system’s performance (baker2018toward, ). However, researchers have begun exploring bi-directional trust in human-robot interactions (rahman2018mutual, ; Wang2014, ).
@@
We grounded our model by confirming that its hidden state is isomorphic to self-reported human trust. Trust model parameters have been estimated using self-reported trust from human participants (chen2020trust, ; akash2017dynamic, ; zahedi2023trust, ; mangalin2024, ). Besides reported trust levels, models have also been estimated using behavioral data, such as action and task performance, from human participants (azevedo2021real, ; akash2020human, ; DBLP:journals/corr/abs-2009-11890, ; williams2023computational, ). Traditionally, when human behavior data is used for model parameter estimation, the state of the system is labeled as the human trust but most work does not directly justify this. In our work, we have verified that our trust model’s estimate is isomorphic to participants’ trust survey responses. With this, we demonstrated that our model accurately captures the dynamics of human trust based solely on behavioral data, without relying on survey responses.
To investigate human trust accurately, a suitable trust-elicitation experiment must be designed to capture trust dynamics effectively. The most prominent experimental paradigm to study human-robot trust relies on block arrangements, wherein the robot is highly reliable in one block and less reliable in the other (salem2015would, ; manzey2012human, ) While these paradigms provide insights into human-robot trust, they do not generate rich enough data to build dynamic models of trust. To address this challenge, we adopted a randomized experimental design wherein the task complexity, influencing robot success rate, is randomly selected in each trial. This approach more naturally excites various levels of human trust and its interaction with different task outcomes, including unexpected and random failures that humans experience when working with robots. Additionally, we believe this experimental paradigm is a closer representation of real-world human-robot interaction scenarios, and is being increasingly adopted in HRI studies (akash2020human, ; azevedo2021real, ; DBLP:journals/corr/abs-2009-11890, )
@@ 7. Conclusions and Future Directions
In the context of human-supervised object collection tasks, we designed an optimal assistance-seeking policy for a mobile manipulator. The human behavior model was formulated as an IOHMM and parameters were estimated using data collected from human subject experiments. The model-based estimates and survey responses were compared and we showed that the estimated trust states from the model are isomorphic to the self-reported trust ratings. We leveraged this model in a POMDP framework to design an optimal assistance-seeking policy that accounts for the human trust state. We showed that the trust-aware policy outperforms a baseline trust-agnostic policy in human-subject experiments. We also showed, by comparing the survey responses of participants, that their trust is higher after working with the robot with a trust-aware policy as compared to working with the robot with a trust-agnostic policy. A possible direction of future research is to schedule low and high-complexity trials for maximizing overall team performance. Other directions of future research include designing trust-aware policies for non-supervisory human-robot collaboration tasks such as active collaboration tasks.

### v2

@@ 6. Discussion
The concept of trust has been explored across various domains, including both human-human and human-robot interactions. Mayer et al. (mayer1995integrative, ) formally define trust between individuals as “the willingness of a party to be vulnerable to the outcomes of another party based on the expectation that the other will perform a particular action important to the trustor, irrespective of the ability to monitor or control that other party.”.
The definition emphasizes that trust becomes crucial when vulnerability, risk, or uncertainty is involved in an interaction. With the deployment of robots working in close proximity with human partners in risky and uncertain environments (takayama2008beyond, ), human trust in robots becomes a crucial factor for their adoption. While human-robot trust shares commonality with human-human trust, the former is typically considered only from the human’s perspective regarding the system’s performance (baker2018toward, ). However, researchers have begun exploring bidirectional trust in human-robot interactions (rahman2018mutual, ; Wang2014, ).
@@
We grounded our model by confirming that its hidden state is isomorphic to self-reported human trust. Trust model parameters have been estimated using self-reported trust from human participants (chen2020trust, ; akash2017dynamic, ; zahedi2023trust, ; mangalin2024, ). Besides reported trust levels, models have also been estimated using behavioral data, such as action and task performance, from human participants (azevedo2021real, ; akash2020human, ; DBLP:journals/corr/abs-2009-11890, ; mangalin2025trustrep, ; williams2023computational, ). Traditionally, when human behavior data is used for model parameter estimation, the state of the system is labeled as the human trust but most work does not directly justify this. In our work, we have verified that our trust model’s estimate is isomorphic to participants’ trust survey responses. With this, we demonstrated that our model accurately captures the dynamics of human trust based solely on behavioral data, without relying on survey responses.
To investigate human trust accurately, a suitable trust-elicitation experiment must be designed to capture trust dynamics effectively. The most prominent experimental paradigm to study human-robot trust relies on block arrangements, wherein the robot is highly reliable in one block and less reliable in the other (salem2015would, ; manzey2012human, ). While these paradigms provide insights into human-robot trust, they do not generate rich enough data to build dynamic models of trust. To address this challenge, we adopted a randomized experimental design wherein the task complexity, influencing robot success rate, is randomly selected in each trial. This approach more naturally excites various levels of human trust and its interaction with different task outcomes, including unexpected and random failures that humans experience when working with robots. Additionally, we believe this experimental paradigm is a closer representation of real-world human-robot interaction scenarios, and is being increasingly adopted in HRI studies (akash2020human, ; azevedo2021real, ; DBLP:journals/corr/abs-2009-11890, )
@@ 7. Conclusions and Future Directions
In the context of human-supervised object collection tasks, we designed an optimal assistance-seeking policy for a mobile manipulator. The human behavior model was formulated as an IOHMM, and parameters were estimated using data collected from human subject experiments. The model-based estimates and survey responses were compared, and we showed that the estimated trust states from the model are isomorphic to the self-reported trust ratings. We leveraged this model in a POMDP framework to design an optimal assistance-seeking policy that accounts for the human trust state. We showed that the trust-aware policy outperforms a baseline trust-agnostic policy in human-subject experiments. We also showed, by comparing the survey responses of participants, that their trust is higher after working with the robot with a trust-aware policy as compared to working with the robot with a trust-agnostic policy. A possible direction of future research is to schedule low and high-complexity trials for maximizing overall team performance. Other directions of future research include designing trust-aware policies for non-supervisory human-robot collaboration tasks such as active collaboration tasks. Additionally, extending to an adaptive human behavior model to model user-specific variations is another direction. This would allow the robot to adapt more effectively by learning and responding to the unique trust dynamics of each user.